OpenStudy (anonymous):

A ball is thrown up with a velocity of 20 m/s. How long will it be before the ball returns to the ground? Assume gravitational acceleration = 10 m/s2. You can ignore air friction.

OpenStudy (anonymous):

What's the weight of the ball?

OpenStudy (anonymous):

You have divide the flight into two parts: First part : The ball is thrown vertically upwards until it has reached its maximum height. Second part: The ball is falling down until it reaches the ground.

OpenStudy (anonymous):

Lets consider the first part: Write the data and use it find the time when it reaches the maximum height: initial velocity= 20m/s acceleration due to gravity= -10m/s2 because it is travelling away from gravity and acceleration is a vector quantity. final velocity= 0m/s because it stops in the air when maximum height is reached. time= ts

OpenStudy (anonymous):

From the equations of motion use the suitable equation to find time which is: v=u+at v-u=at v-u/a=t 0-20/-10=t t=2 seconds

OpenStudy (anonymous):

The ball took 2 seconds to reach the maximum height and will take the same time to return to the ground so: 2+2= 4 seconds

OpenStudy (anonymous):

The ball is in the air for 4 seconds.

OpenStudy (anonymous):

formula is 2u/g therefore-2*20/10=4sec.

OpenStudy (anonymous):

Unless you know the weight of the ball and the speed at which it left the hand. The speed of the ball is not a constant if gravity is involved. Air friction would play a part on size, but the weight still is a factor. If the speed is 20 m/s at start then weight has to be a factor. Question isn't broken down enough.